# A method for sparse and robust independent component analysis

### v2

@@ Abstract
This work presents sparse invariant coordinate selection, SICS, a new method for sparse and robust independent component analysis. SICS is based on classical invariant coordinate selection, which is presented in such a form that a LASSO-type penalty can be applied to promote sparsity. Robustness is achieved by using robust scatter matrices. In the first part of the paper, the background and building blocks: scatter matrices, measures of robustness, ICS and independent component analysis, are carefully introduced. Then the proposed new method and its algorithm are derived and presented. This part also includes a consistency result for a general case of sparse ICS-like methods. The performance of SICS in identifying sparse independent component loadings is investigated with multiple simulations. The method is illustrated with an example in constructing sparse causal graphs and we also propose a graphical tool for selecting the appropriate sparsity level in SICS.
@@ 1 Introduction
The paper is organized as follows. In Section 2, we review the concepts of a scatter matrix and independence property, and recall how they produce a solution to the ICA problem. Also several robust scatter matrices from the literature are presented. Section 3 sees us combining robust scatter matrices with the sparsity framework of [31] to obtain our proposed method. We also establish the convergence rate of the method in this section. In Section 4, we compare the performance of the method under different levels of sparsity and robustness using simulations, including also several competing methods in the evaluations. A real data example of estimating non-Gaussian causal graphs using sparse ICA is given in Section 5. A graphical tool for selecting the sparsity level is proposed in Section 6, and we conclude with discussion in Section 7.
@@ 2.1 ICA solution using two scatter matrices
Now, a question rises, how can we find scatter matrices with the independence property? Besides the covariance matrix, another well-known example is the FOBI-matrix (used in an ICA method named fourth order blind identification [5]) $S_{\mathrm{FOBI}}=\mathbb{E}[\tilde{x}\tilde{x}^{\prime}\Sigma^{-1}\tilde{x}\tilde{x}^{\prime}$], where $\tilde{x}:=x-\mathbb{E}(x)$ and $\Sigma$ denotes the covariance matrix of $x$. Moreover, further scatter matrices with the independence property can be obtained via the process call symmetrization, and we next state [44, Theorem 1] defining symmetrized scatter matrices
@@
[44] Let $S$ be a scatter matrix and $x_{1}$, $x_{2}$ two independent copies of a random vector $x$. Then the symmetrized scatter matrix
$S_{s}(x):=S(x_{1}-x_{2})$
has the independence property.
@@
In Theorem 2, the location vector $T$ is used simply to fix the location of the independent components. Even without $T$, the ICs could be estimated up to location as $B(x)^{\prime}x$ and, as such, we focus solely on the scatter matrices $S_{1}$ and $S_{2}$ in the following, disregarding the location estimation. Independent component analysis based on Theorem 2 using symmetrized scatter matrices is discussed in [52]. Finally, we note that Theorem 2 indeed states that any two scatter matrices with the independence property can be used to solve the IC problem on the population level. However, these scatter matrices might involve various assumptions and their finite-sample properties can still be different. Later we focus on a particular class of scatters that allows us to solve the problem in an outlier-resistant way.
@@ 2.2 Regression formulation for scatter matrix ICA
$\sum_{j=1}^{p}\|S_{1}(x)^{-1/2}r_{j}-AB^{\prime}r_{j}\|_{2}^{2}$
over $B\in\mathbb{R}^{p\times k}, AA$$^{\prime}A=I_{k}$, are precisely the pairs
$(A,B)=(U_{k}O_{k},S_{1}(x)^{-1/2}U_{k}O_{k}),$
where $O_{k}$ is any $k\times k$ orthogonal matrix.
@@
$-2\mathrm{tr}(B^{\prime}S_{2}(x)S_{1}^{-1/2}(x)A)+\mathrm{tr}(B^{\prime}S_{2}(x)B).$
Differentiating this with respect to $B$, we get the gradient
$-2S_{2}(x)S_{1}^{-1/2}A+2S_{2}B.$
The positive definiteness of $S_{2}$ thus shows that the minimizing value of $B$ satisfies $B=S_{1}^{-1/2}(x)A$. Plugging this back in to the objective function, we see that the optimal $A$ maximizes the map
$A\mapsto\mathrm{tr}(A^{\prime}S_{1}^{-1/2}(x)S_{2}(x)S_{1}^{-1/2}(x)A),$
over $\in\mathbb{R}^{p\times k}, AA$$^{\prime}A=I_{k}$. By [23, Corollary 4.3.39], all maximizers are thus of the form $A=V_{k}O_{k}$ where $V_{k}\in\mathbb{R}^{p\times k}$ contain first $k$ eigenvectors of $S_{1}^{-1/2}(x)S_{2}(x)S_{1}^{-1/2}(x)$ as its columns and $O_{k}$ is an arbitrary $k\times k$ orthogonal matrix. By the affine equivariance of scatter matrices, $U_{2}\left(S_{1}(x)^{-1/2}x\right)$ in Theorem 2 contains the $p$ eigenvectors of $S_{2}(S_{1}^{-1/2}(x)x)=S_{1}^{-1/2}(x)S_{2}(x)S_{1}^{-1/2}(x)$, showing that $V_{k}=U_{k}$ (up to sign), and concluding the proof. ∎
@@ 2.3 Symmetrized robust scatter matrices
$L(S)=\frac{1}{n}\sum_{i=1}^{n}\{\rho(x_{i}^{\prime}S^{-1}x_{i})-\rho(x_{i}^{\prime}x_{i})\}+\ln\det S.$
This expression originally comes from the maximum likelihood estimation of an elliptic distribution with density proportional to $\mathrm{det}(S^{-1})\exp(-\rho(x^{\prime}S^{-1}x))$. The function $\rho$ is typically assumed to satisfy various regularity conditions, see [16, 14]. One common choice for the function $\rho$ is $\rho_{\nu,p}(z)=(\nu+p)\ln(\nu+z)$ which corresponds to the density of the $p$-variate $t$-distribution with $\nu>0$ degrees of freedom.
@@
$\frac{1}{n}\sum_{i=1}^{n}w\left(\|S^{-1/2}x_{i}\|^{2}\right)\|S^{-1/2}x_{i}\|^{-2}S^{-1/2}x_{i}x_{i}^{\prime}S^{-1/2}=I_{p},$ (1)
where $w(z)=\rho^{\prime}(z)z$ is called a weight function. $M$-estimators are often defined in terms of the function $w$ instead of the function $\rho$. Common choices for $w$ are [49] $w(z)=z$ giving the regular covariance matrix and $w(z)=p$ giving Tyler’s $M$-estimator [53]. Another common choice is the Huber’s $M$-estimator
$w(z)=\begin{cases}|z|/\sigma,\ &\text{for }|z|\leq c\\ c/\sigma,\ &\text{for }|z|>c,\end{cases}$
where the tuning constant $c$ is chosen to control $\mathbb{P}(\chi^{2}_{p}\leq c^{2}/2)$ and the scaling factor $\sigma$ is chosen so that $\mathbb{E}[w(\|x\|^{2})]=p$ for $x\sim N(0,I_{p})$.
@@
Naturally, other choices of symmetrized scatters could also be equally viable in our context. Scatter matrices can be categorized into three groups [55]: Class I contains all non-robust scatters (such as $\mathrm{Cov}(x)$ and $S_{\mathrm{FOBI}}$), Class II contains scatters that are moderately robust and have relatively light computational complexity (such as $M$-estimators), and Class III contains scatters with high robustness and high computational complexity (such as the minimum covariance determinant matrix [47]). [55, Section 6] state that using two Class II matrices is typically a reasonable choice in ICS when one wants to avoid the computational load of Class III scatters, which in our case would still be increased due to symmetrization. And, while class II contains also other matrices besides our pair of choice, such as Dümbgen’s estimator [12] or one-step $W$-estimators [55], our primary focus is on sparsity and thus we have not attempted to perform any extensive evaluations between different robust scatters here. However, several comparisons of symmetrized robust scatter matrices can be found in the literature, see [49, 52, 37].
@@ 2.4 Measuring robustness
The two most common ways to measure robustness of a statistic are the breakdown point and influence function. Breakdown point describes how large proportion of observations can be outliers without the statistic having arbitrary large/small values. Let $T$ be a statistic taking values in some normed space, $X$ a sample and $X_{\varepsilon}$ a sample with proportion $\varepsilon$ of the observation replaced with arbitrary values. Let
$b(X_{\varepsilon},T)=\sup_{X_{\varepsilon}}\|T(X)-T(X_{\varepsilon})\|$
@@ 3.1 Algorithm
Based on the concepts in the previous section, we next propose an algorithm for obtaining a robust and sparse ICA solution. Let $S_{1}$ and $S_{2}$ be two robust scatter matrices with the independence property, $r_{j}\in\mathbb{R}^{p}$ denote the columns of $S_{2}(x)^{1/2}$ and $\lambda_{m}\geq 0$ be the sparsity penalization parameter corresponding to the $m$th independent component. Then, the matrix $B=(\beta_{1},\dots,\beta_{k})$ part of the minimizer $A,B\in\mathbb{R}^{p\times k}$ $(A^{\prime}A~=~I_{k})$ of
$\sum_{j=1}^{p}\|S_{1}(x)^{-1/2}r_{j}-AB^{\prime}r_{j}\|_{2}^{2}+\sum_{m=1}^{k}\lambda_{m}\|\beta_{m}\|_{1}$ (2)
@@ 5 $B_{temp}=B$;
7    $\beta_{j}=\arg\min_{\beta_{j}}\left\{\sum\|S_{2}^{\frac{1}{2}}S_{1}^{-\frac{1}{2}}\alpha_{j}-S_{2}^{\frac{1}{2}}\beta_{j}\|_{2}^{2}+\lambda_{j}\|\beta_{j}\|_{1}\right\}$;
@@ 9 $B=\text{fix\_signs}(B)$;
10   Calculate the singular value decomposition $S_{1}^{-\frac{1}{2}}S_{2}B=UDV^{\prime}$;
11 $A_{0}=UV^{\prime}$;
12   Calculate the eigendecomposition $A_{0}^{\prime}S_{1}^{-\frac{1}{2}}S_{2}S_{1}^{-\frac{1}{2}}A_{0}=O_{k}\Delta O_{k}^{\prime}$;
@@ 13 $A=A_{0}O_{k}$;
$\operatorname{tr}((I_{p}-AA^{\prime})S_{1}^{-1}S_{2})+\|S_{2}^{\frac{1}{2}}S_{1}^{-\frac{1}{2}}A-S_{2}^{\frac{1}{2}}B\|^{2}+\sum_{m=1}^{k}\lambda_{m}\|\beta_{m}\|_{1},$
@@
$\sum_{j=1}^{p}\|S_{1}^{-1/2}r_{j}-AB^{\prime}r_{j}\|_{2}^{2}=\|S_{1}^{-1/2}S_{2}^{1/2}-AB^{\prime}S_{2}^{1/2}\|_{2}^{2}=\|S_{2}^{1/2}S_{1}^{-1/2}-S_{2}^{1/2}S_{1}^{-1/2}S_{1}^{1/2}BA^{\prime}\|_{2}^{2}$
and using the reduced rank Procrustes rotation [61, Theorem 4] to minimize this expression with respect to $A$. As this estimate is only calculated up to the rotation $O_{k}$, we choose $O_{k}$ (lines 12–13) so that $A$ diagonalizes $S_{1}^{-\frac{1}{2}}S_{2}S_{1}^{-\frac{1}{2}}$. This choice thus corresponds to how the solution to the joint diagonalization behaves in the regular ICS.
Algorithm 1 thus produces a sparse matrix $B$ (whose sparsity level is controlled by the parameters $r_{1},\ldots,r_{k}$), using which the independent components are obtained as $B^{\prime}x_{i}$.

### v3

@@ Abstract
This work presents sparse invariant coordinate selection, SICS, a new method for sparse and robust independent component analysis. SICS is based on classical invariant coordinate selection, which is presented in such a form that a LASSO-type penalty can be applied to promote sparsity. Robustness is achieved by using robust scatter matrices. In the first part of the paper, the background and building blocks: scatter matrices, measures of robustness, ICS and independent component analysis, are carefully introduced. Then the proposed new method and its algorithm are derived and presented. This part also includes consistency and breakdown point results for a general case of sparse ICS-like methods. The performance of SICS in identifying sparse independent component loadings is investigated with multiple simulations. The method is illustrated with an example in constructing sparse causal graphs and we also propose a graphical tool for selecting the appropriate sparsity level in SICS.
@@ 1 Introduction
    We characterize the breakdown point of SICS in terms of the breakdown points of the two scatter matrices.
  • 4.
@@
In Section 2, we review the concepts of a scatter matrix and independence property, and recall how they produce a solution to the ICA problem. Also several robust scatter matrices from the literature are presented. Section 3 sees us combining robust scatter matrices with the sparsity framework of [31] to obtain our proposed method. We also establish the convergence rate and the breakdown point of the method in this section. In Section 4, we compare the performance of the method under different levels of sparsity and robustness using simulations, including also several competing methods in the evaluations. A real data example of estimating non-Gaussian causal graphs using sparse ICA is given in Section 5. A graphical tool for selecting the sparsity level is proposed in Section 6, and we conclude with discussion in Section 7. The following notation will be used for different norms: $\|\cdot\|$ denotes the Euclidean vector norm, $\|\cdot\|_{1}$ the $L_{1}$-norm for vectors, $\|\cdot\|_{F}$ the Frobenius norm for matrices and $\|\cdot\|_{2}$ the spectral norm for matrices.
@@ 2.1 ICA solution using two scatter matrices
Now, a question rises, how can we find scatter matrices with the independence property? Besides the covariance matrix, another well-known example is the FOBI-matrix (used in an ICA method named fourth order blind identification [5]) $S_{\mathrm{FOBI}}=\mathbb{E}[\tilde{x}\tilde{x}^{\top}\Sigma^{-1}\tilde{x}\tilde{x}^{\top}$], where $\tilde{x}:=x-\mathbb{E}(x)$ and $\Sigma$ denotes the covariance matrix of $x$. Moreover, further scatter matrices with the independence property can be obtained via the process call symmetrization, and we next state [44, Theorem 1] defining symmetrized scatter matrices
@@
[44] Let $S$ be a scatter matrix and $x_{1}$, $x_{2}$ two independent copies of a random vector $x$. Then the symmetrized scatter matrix $S_{s}(x):=S(x_{1}-x_{2})$ has the independence property.
@@
In Theorem 2, the location vector $T$ is used simply to fix the location of the independent components. Even without $T$, the ICs could be estimated up to location as $B(x)^{\top}x$ and, as such, we focus solely on the scatter matrices $S_{1}$ and $S_{2}$ in the following, disregarding the location estimation. Independent component analysis based on Theorem 2 using symmetrized scatter matrices is discussed in [53]. Finally, we note that Theorem 2 indeed states that any two scatter matrices with the independence property can be used to solve the IC problem on the population level. However, these scatter matrices might involve various assumptions and their finite-sample properties can still be different. Later we focus on a particular class of scatters that allows us to solve the problem in an outlier-resistant way.
@@ 2.2 Regression formulation for scatter matrix ICA
$\sum_{j=1}^{p}\|S_{1}(x)^{-1/2}r_{j}-AB^{\top}r_{j}\|^{2}$
over $B\in\mathbb{R}^{p\times k}, AA$$^{\top}A=I_{k}$, are precisely the pairs $(A,B)=(U_{k}O_{k},S_{1}(x)^{-1/2}U_{k}O_{k})$, where $O_{k}$ is any $k\times k$ orthogonal matrix.
@@
$-2\mathrm{tr}(B^{\top}S_{2}(x)S_{1}^{-1/2}(x)A)+\mathrm{tr}(B^{\top}S_{2}(x)B).$
Differentiating this with respect to $B$, we get the gradient $2S_{2}(x)S_{1}^{-1/2}A+2S_{2}B$. The positive definiteness of $S_{2}$ thus shows that the minimizing value of $B$ satisfies $B=S_{1}^{-1/2}(x)A$. Plugging this back in to the objective function, we see that the optimal $A$ maximizes the map $A\mapsto\mathrm{tr}(A^{\top}S_{1}^{-1/2}(x)S_{2}(x)S_{1}^{-1/2}(x)A)$ over $\in\mathbb{R}^{p\times k}, AA$$^{\top}A=I_{k}$. By [23, Corollary 4.3.39], all maximizers are thus of the form $A=V_{k}O_{k}$ where $V_{k}\in\mathbb{R}^{p\times k}$ contain first $k$ eigenvectors of $S_{1}^{-1/2}(x)S_{2}(x)S_{1}^{-1/2}(x)$ as its columns and $O_{k}$ is an arbitrary $k\times k$ orthogonal matrix. By the affine equivariance of scatter matrices, $U_{2}\left(S_{1}(x)^{-1/2}x\right)$ in Theorem 2 contains the $p$ eigenvectors of $S_{2}(S_{1}^{-1/2}(x)x)=S_{1}^{-1/2}(x)S_{2}(x)S_{1}^{-1/2}(x)$, showing that $V_{k}=U_{k}$ (up to sign), and concluding the proof. ∎
@@ 2.3 Symmetrized robust scatter matrices
$L(S)=\frac{1}{n}\sum_{i=1}^{n}\{\rho(x_{i}^{\top}S^{-1}x_{i})-\rho(x_{i}^{\top}x_{i})\}+\ln\det S.$
This expression originally comes from the maximum likelihood estimation of an elliptic distribution with density proportional to $\mathrm{det}(S^{-1})\exp(-\rho(x^{\top}S^{-1}x))$. The function $\rho$ is typically assumed to satisfy various regularity conditions, see [16, 14]. One common choice for the function $\rho$ is $\rho_{\nu,p}(z)=(\nu+p)\ln(\nu+z)$ which corresponds to the density of the $p$-variate $t$-distribution with $\nu>0$ degrees of freedom.
@@
$\frac{1}{n}\sum_{i=1}^{n}w\left(\|S^{-1/2}x_{i}\|^{2}\right)\|S^{-1/2}x_{i}\|^{-2}S^{-1/2}x_{i}x_{i}^{\top}S^{-1/2}=I_{p},$ (1)
where $w(z)=\rho^{\prime}(z)z$ is called a weight function. $M$-estimators are often defined in terms of the function $w$ instead of $\rho$. Common choices for $w$ are [50] $w(z)=z$ giving the regular covariance matrix, $w(z)=p$ giving Tyler’s $M$-estimator [54], and Huber’s $M$-estimator which uses $w(z)=\mathbb{I}(|z|\leq c)|z|/\sigma+\mathbb{I}(|z|>c)c/\sigma$, where the tuning constant $c$ is chosen to control $\mathbb{P}(\chi^{2}_{p}\leq c^{2}/2)$ and the scaling factor $\sigma$ is chosen so that $\mathbb{E}[w(\|x\|^{2})]=p$ for $x\sim N(0,I_{p})$.
@@
Naturally, other choices of symmetrized scatters could also be equally viable in our context. Scatter matrices can be categorized into three groups [56]: Class I contains all non-robust scatters (such as $\mathrm{Cov}(x)$ and $S_{\mathrm{FOBI}}$), Class II contains scatters that are moderately robust and have relatively light computational complexity (such as $M$-estimators), and Class III contains scatters with high robustness and high computational complexity (such as the minimum covariance determinant matrix [47]). [56, Section 6] state that using two Class II matrices is typically a reasonable choice in ICS when one wants to avoid the computational load of Class III scatters, which in our case would still be increased due to symmetrization. And, while class II contains also other matrices besides our pair of choice, such as Dümbgen’s estimator [12] or one-step $W$-estimators [56], our primary focus is on sparsity and thus we have not attempted to perform any extensive evaluations between different robust scatters here. However, several comparisons of symmetrized robust scatter matrices can be found in the literature, see [50, 53, 37].
@@ 2.4 Measuring robustness
The two most common ways to measure robustness of a statistic are the breakdown point and influence function. Breakdown point describes how large proportion of observations can be outliers without the statistic having arbitrary large/small values. Let $T$ be a statistic taking values in some normed space $(\mathcal{T},\|\cdot\|_{\mathcal{T}})$, $X$ a sample and $X_{\varepsilon}$ a sample with proportion $\varepsilon$ of the observation replaced with arbitrary values, i.e., an $\varepsilon$-contamination of $X$. Let
$b(X_{\varepsilon},T)=\sup_{X_{\varepsilon}}\|T(X)-T(X_{\varepsilon})\|_{\mathcal{T}}$
@@ 3.1 Algorithm
Based on the concepts in the previous section, we next propose an algorithm for obtaining a robust and sparse ICA solution. Let $S_{1}$ and $S_{2}$ be two robust scatter matrices with the independence property, $r_{j}\in\mathbb{R}^{p}$ denote the columns of $S_{2}(x)^{1/2}$ and $\lambda_{m}\geq 0$ be the sparsity penalization parameter corresponding to the $m$th independent component. Then, the matrix $B=(\beta_{1},\dots,\beta_{k})$ part of the minimizer $A,B\in\mathbb{R}^{p\times k}$ $(A^{\top}A~=~I_{k})$ of
$\sum_{j=1}^{p}\|S_{1}(x)^{-1/2}r_{j}-AB^{\top}r_{j}\|^{2}+\sum_{m=1}^{k}\lambda_{m}\|\beta_{m}\|_{1}$ (2)
@@ 5 $B_{temp}=B$;
7    $\beta_{j}=\arg\min_{\beta_{j}}\left\{\sum\|S_{2}^{\frac{1}{2}}S_{1}^{-\frac{1}{2}}\alpha_{j}-S_{2}^{\frac{1}{2}}\beta_{j}\|^{2}+\lambda_{j}\|\beta_{j}\|_{1}\right\}$;
@@ 9 $B=\text{fix\_signs}(B)$;
10   Calculate the singular value decomposition $S_{1}^{-\frac{1}{2}}S_{2}B=UDV^{\top}$;
11 $A_{0}=UV^{\top}$;
12   Calculate the eigendecomposition $A_{0}^{\top}S_{1}^{-\frac{1}{2}}S_{2}S_{1}^{-\frac{1}{2}}A_{0}=O_{k}\Delta O_{k}^{\top}$;
@@ 13 $A=A_{0}O_{k}$;
$\operatorname{tr}((I_{p}-AA^{\top})S_{1}^{-1}S_{2})+\|S_{2}^{\frac{1}{2}}S_{1}^{-\frac{1}{2}}A-S_{2}^{\frac{1}{2}}B\|_{F}^{2}+\sum_{m=1}^{k}\lambda_{m}\|\beta_{m}\|_{1},$
@@
$\sum_{j=1}^{p}\|S_{1}^{-1/2}r_{j}-AB^{\top}r_{j}\|^{2}=\|S_{1}^{-1/2}S_{2}^{1/2}-AB^{\top}S_{2}^{1/2}\|_{F}^{2}=\|S_{2}^{1/2}S_{1}^{-1/2}-S_{2}^{1/2}S_{1}^{-1/2}S_{1}^{1/2}BA^{\top}\|_{F}^{2}$
and using the reduced rank Procrustes rotation [62, Theorem 4] to minimize this expression with respect to $A$. As this estimate is only calculated up to the rotation $O_{k}$, we choose $O_{k}$ (lines 12–13) so that $A$ diagonalizes $S_{1}^{-\frac{1}{2}}S_{2}S_{1}^{-\frac{1}{2}}$. This choice thus corresponds to how the solution to the joint diagonalization behaves in the regular ICS.
Algorithm 1 thus produces a sparse matrix $B$ (whose sparsity level is controlled by the parameters $r_{1},\ldots,r_{k}$), using which the independent components are obtained as $B^{\top}x_{i}$.
@@ 3.2 Consistency of SICS
3.3 Robustness of SICS
In this section, we give a result showing that, when using robust scatter matrices, the SICS algorithm gives robust results. It is clear that when a scatter matrix breaks in sense of breakdown point, algorithms based on it also break. The following Theorem 5 states that if neither of the scatter matrices break, also SICS does not break. Therefore, the breakdown point of SICS satisfies $\varepsilon^{*}(\text{SICS})=\min(\varepsilon^{*}(S_{1}),\varepsilon^{*}(S_{2}))$. The proof of the theorem is given in C. Let us denote in the following the set of $p\times p$ positive definite matrices by $\mathcal{S}_{+}^{p}$ and the $k$th largest eigenvalue of a matrix $S$ by $\phi_{k}(S)$.
Theorem 5.
Let $f(A,B;S_{1},S_{2})=\|S_{1}^{-\frac{1}{2}}S_{2}^{\frac{1}{2}}-AB^{\top}S_{2}^{\frac{1}{2}}\|_{F}^{2}+\lambda\sum_{m=1}^{k}\|\beta_{m}\|_{1}$ be the SICS objective function with $B=(\beta_{1},\dots,\beta_{k})\in\mathbb{R}^{p\times k}, AA$$^{\top}A~=~I_{k}$ and fixed $\lambda\geq 0$. Let $(S_{10},S_{20}):=(S_{1}(X),S_{2}(X))$ be the scatters $S_{1},S_{2}$ evaluated for some fixed sample $X$. Fix $\varepsilon>0$ and let
$\mathcal{G}=\{(S_{1}(X_{\varepsilon}),S_{2}(X_{\varepsilon}))\mid X_{\varepsilon}\mbox{ is an }\varepsilon\mbox{-contamination of }X\}\subset\mathcal{S}_{+}^{p}\times\mathcal{S}_{+}^{p}.$
Assume that the following hold:
  • (i)
    We have $\inf_{(S_{1},S_{2})\in\mathcal{G}}\phi_{p}(S_{2})>0$.
  • (ii)
    We have $\sup_{(S_{1},S_{2})\in\mathcal{G}}\|S_{1}^{-\frac{1}{2}}-S_{10}^{-\frac{1}{2}}\|_{F}<\infty$ and $\sup_{(S_{1},S_{2})\in\mathcal{G}}\|S_{2}-S_{20}\|_{F}<\infty$.
Then, letting $(A(S_{1},S_{2}),B(S_{1},S_{2}))$ denote any minimizer of $(A,B)\mapsto f(A,B;S_{1},S_{2})$, we have
$\sup_{(S_{1},S_{2})\in\mathcal{G}}\|B(S_{1},S_{2})-B(S_{10},S_{20})\|_{F}<\infty.$
Theorem 5 is stated in terms of samples $X$, but it extends trivially to general distributions and their contaminations. The conditions in assumption $(ii)$ in Theorem 5 are asymmetric in $S_{1}$ and $S_{2}$ due to the form of the objective function. However, Theorem 9 in C gives that the first condition in assumption $(ii)$ holds if we assume $\sup_{(S_{1},S_{2})\in\mathcal{G}}\|S_{1}-S_{10}\|_{F}<\infty$ and $\inf_{(S_{1},S_{2})\in\mathcal{G}}\phi_{p}(S_{1})>0$, leading to sufficient conditions that are symmetric in $S_{1},S_{2}$. Finally, we note that by taking $\lambda=0$, Theorem 5 covers also the robustness of regular ICS.